xx TestComplex

TestComplex

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11imaginaryxvalue 10Scope: privatex01  REAL

11Description: power of the complex number to take

11 Range 10 x01 Default: -2
11*:*  10imaginary part

11

11powerxvalue 10Scope: privatex01  REAL

11Description: power of the complex number to take

11 Range 10 x01 Default: 0.5
110.0:  10Some positive value

11

11realxvalue  10Scope: privatex01  REAL

11Description: power of the complex number to take

11 Range 10 x01 Default: 3
11*:*  10real part

11

7 Interfaces

General Implements: testcomplex

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11  cvars3d  10  x01  compact110
11                 10  comp3d  x01 description113D complex grid functions
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels112
11                 10  x01variable type11COMPLEX
11  carrays3d  10  x01  compact110
11                 10  acomp3d  x01 description113D complex grid arrays
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110
11                 10  x  x01  size1110
11                 10  x01  size1110

11                 10  x01 timelevels111
11                 10  x01variable type11COMPLEX
11  carrays2d  10  x01  compact110
11                 10  acomp2d  x01 description112D complex grid arrays
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110
11                 10  x  x01  size1110

11                 10  x01 timelevels111
11                 10  x01variable type11COMPLEX
11  carrays1d  10  x01  compact110
11                 10  acomp1d  x01 description111D complex grid arrays
11                 10  x01 dimensions111
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110

11                 10  x01 timelevels111
11                 10  x01variable type11COMPLEX
11  cscalars  10  x01  compact110
11                 10  scomp  x01 description11complex grid scalars
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11COMPLEX
11 cmplxxoutput 10  x01  compact110
11                 10  realxpart  x01 dimensions110
11                 10 imaginaryxpart x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11REAL

11

8 Schedule This section lists all the variables which are assigned storage by thorn CactusTest/TestComplex. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage

11Always:  11x
11cvars3d[2]  11x
11carrays3d carrays2d carrays1d cscalars11x
11cmplxxoutput  11x

11x  11x

11

Scheduled Functions

CCTKxINITIAL x initcomplex3d x initialise 3d complex grid functions x

11x11Language:11fortran
11 x 11Type: 11function

11

CCTKxEVOL x evolvecomplex3d x evolve 3d complex grid functions x

11x11Language:11fortran
11 x 11Sync:  11cvars3d
11 x 11Type: 11function

11

CCTKxPOSTINITIAL x testcomplexpower x do a simple complex power operation x

11x11Language:11c
11 x 11Type: 11function